a 15 gram gun fires a bullet a a muzzle velocuty of 600 m/s. What is the recoil velocity of the gun?

1
Expert's answer
2022-02-19T04:56:06-0500

Given:

"m_{gun}=15\:\rm g"

"m_{bullet}=2\:\rm g"

"v_{bullet}=600\:\rm m/s"


The law of conservation of momentum says

"m_{bullet}v_{bullet}=m_{gun}v_{gun}"

Hence, the recoil speed of the gun

"v_{gun}=\frac{m_{bullet}v_{bullet}}{m_{gun}}=\frac{2*600}{15}=80\:\rm m/s"